OpenStudy (lollola):

determine whether the sequence 16,8,4,2 is arithmetic, geometric, both, or neither

OpenStudy (anonymous):

it's geometric sequence if I made the correct calculations

OpenStudy (lollola):

Ok thanks

OpenStudy (anonymous):

this is how I decided : 16;8;4;2 First we check for arithmetic sequence,we know that arithmetical sequence works like this : a2-a1=a4-a3 , so 8-16=2-4 ( which isn't true -8=-2) So it isn't arithmetic. Now for geomethric : 8/16=2/4 ( which is true we get 1/2=1/2)